9. A dehydrated patient needs a 6.4% saline IV. Unfortunately, the hospital only has bags of 6% and 8% saline solutions. How many liters of each of these solutions should be mixed together to yield 1 liter of the desired concentration?

Hello, can someone, anyone please help me with this, thanks so much!!

Answer by stanbon(75887)

A dehydrated patient needs a 6.4% saline IV. Unfortunately, the hospital only has bags of 6% and 8% saline solutions. How many liters of each of these solutions should be mixed together to yield 1 liter of the desired concentration?
--------------------
Equation:
salt + salt = salt
----
0.06x + 0.08(1-x) = 0.064*1
-----
Multiply thru by 1000 to get:
60x + 80 - 80x = 64
------------------------
-20x = -16
x = 0.8 liters (amt. of 6% solution needed)
1-x = 0.2 liters (amt. of 8% solution needed)
================================================
Cheers,
Stan H.